OpenStudy (anonymous):

Already Answered! What is the equation of the graph? It's a graph with a circle on it..I NEED SOMEONE TO check, please!?

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

(x-3)^2+(y-(-7))^2=(3)^2=9 (x-3)^2+(y+7)^2=9

OpenStudy (anonymous):

;) keep rocking

OpenStudy (anonymous):

yeah, i was wrong, it's (y+7)^2